OpenStudy (anonymous):

can anyone show me how to get the solution to this radical equation x-3+4=6

OpenStudy (anonymous):

\[\sqrt{x-3}+4=6\] maybe?

OpenStudy (anonymous):

you wrote \[x-3+4=6\] which has no radical, so i am guessing

OpenStudy (anonymous):

@ satellite73 yes that's exactly it the first way you made it

OpenStudy (anonymous):

ok take three steps 1) subtract 4 from both sides 2) square both sides to get rid of the radical on the left 3) add 3 to both sides to get \(x\) by itself

OpenStudy (anonymous):

@ satellite73 im still confuse is it 1

OpenStudy (anonymous):

\[\sqrt{x-3}+4=6\] subtract 4 from both sides to get \[\sqrt{x-3}=2\] then square both sides

OpenStudy (anonymous):

you get \[x-3=7\] now solve for \(x\) by adding 3 to both sides

OpenStudy (anonymous):

@ satellite73 i added 3 on both sides i got 6 and 10 what do i do next